OpenStudy (anonymous):

Use conditions Slope = -2, passing through (0, -3) to write an equation for each line in point-slope form & slope-intercept form

OpenStudy (ibbutibbu.):

y = mx + b -3 = -2(0) + b -3=b y = -2x - 3